Grade 12th passAlgebra

Find five rational numbers between.
2/3 and 4/5.........

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us make the denominators same, say 60
i.e., 2/3 and 4/5 can be written as:
2/3=(2×20)/(3×20)=40/60
4/5=(4×12)/(5×12)=48/60
Five rational numbers between 2/3 and 4/5 = five rational numbers between 40/60 and 48/60
Therefore, Five rational numbers between 40/60 and 48/60 = 41/60, 42/60, 43/60, 44/60, 45/60

Thanks